What is the(square root 3 + 2)(sqaure root 6-5)
Answer by dabanfield(803)

What is the(square root 3 + 2)(sqaure root 6-5)
sqrt(3+2) = sqrt(5)
sqrt(6-5) = sqrt(1) = 1
sqrt(3+2)*sqrt(6-5) = sqrt(5)* 1 = sqrt(5)